# 4K4DGen: Panoramic 4D Generation at 4K Resolution

### v2

@@ 4.2 Results
Comparison to Baselines.
We present both qualitative and quantitative comparisons of 4K4DGen to 3D-Cinemagraphy in Fig. 5 and Tab. 1. The figure illustrates that 4K4DGen generates results of superior quality. According to the data presented in the table, 4K4DGen not only achieves higher CLIP similarity scores but also receives more favorable evaluations from users regarding its quality.
Visualization Results.
Additional visualization results, encompassing both rendered images and depth maps, are displayed in Figure 4. These qualitative results demonstrate that 4K4DGen can effectively generate 4D panoramic scenes that are both spatially and temporally consistent. This capability significantly enhances the immersive experience on VR devices.
Figure 6: Comparison to Different Animators: An animator primarily trained on perspective images, when applied to panoramas, is likely to generate minor motion, and the resolution may be limited. Conversely, animating perspective images can result in inconsistencies between different views.
@@ 4.3 Ablation Studies
We perform ablation studies for both the animating and lifting phases of our methodology. During the animating phase, we demonstrate the critical role of our spherical denoise strategy by substituting it with two basic animation strategies. In the lifting phase, we examine the effects of omitting the Spatial-Temporal Geometry Alignment process and the temporal loss in optimizing 4D representations.
@@
To animate the panorama into a panoramic video, a naive approach is to apply animators directly to the entire panorama. However, we observed that this strategy often results in minor motion, as shown in Fig. 6 (b) and Tab. 2 (Animate Pano.). This issue arises due to two main reasons: (1) animators are typically trained on perspective images with a narrow field of view (FoV), whereas panoramas have a 360-degree FoV with specific distortions under the equirectangular projection; (2) our panorama is high-resolution (4K), which exceeds the training distribution of most 2D animators and can easily cause out-of-memory issues, even with an 80GB VRAM graphics card. Thus we have to resize the panorama to a small resolution (2k), causing a loss of details. Thus, we seek to animate on perspective views. Applying the animator to perspective views offers benefits such as reduced distortion and domain-appropriate input for the animator, allowing for smooth animation of high-resolution panoramas. However, animating perspective images separately can introduce inconsistencies between overlapping perspective views, as illustrated in Fig. 6 (c) and Tab. 2 (Animate Pers.). To resolve this challenge, we propose simultaneously denoising all perspective views and fusing them at each denoising step, which capitalizes on the benefits of animating perspective views while ensuring cross-view consistency. The results are displayed in Fig. 6 (a) and Tab. 2 (Ours).
@@
Table 2: Different Animation Strategies in the Animating Phase. Animating the entire panorama results in minor motion and reduced resolution (first row). Conversely, animating from perspective views leads to inconsistencies across different views (second row). This is supported by user feedback, where 70% of participants identified our method as view-consistent, in contrast to only 33% who found animation from perspective views to be view-consistent.
@@
Animate Pano. $2048\times 1024$ 0.983 0.30 -
Animate Pers. $4096\times 2048$ 0.985 0.91 33%
Ours $4096\times 2048$ 0.979 1.27 70%
5 Conclusion, Limitation and Broader Impact
@@ Conclusion.
Broader Impact.

### v3

@@ 4.2 Results
Quantitative Results.
We also show the qualitative comparison between 4K4DGen and 3D-Cinemagraphy (Li et al., 2023a) in Tab. 1, considering both frame and video quality. In terms of frame quality, 4K4DGen achieves significantly better performance on both distribution-based metrics and LLM-based metrics. In terms of video quality, 4K4DGen achieves better Q-Align score, and is selected as the method with best visual quality by 81% of users in our study.
Qualitative Results.
We present a qualitative comparison between 4K4DGen and 3D-Cinemagraphy (3D-Cin.). Since the performance of 3D-Cin. is similar under the "circle" and "zoomin" settings in Tab. 1, we use the "circle" setting to represent 3D-Cin. in Fig. 4. As shown in the figure, 4K4DGen produces high-quality perspective videos that maintain consistency across both time and views, whereas 3D-Cin. struggles with generating ghosting artifacts in the middle frames.
Figure 5: Comparison to Different Animators: Animators trained primarily on perspective images tend to produce limited motion when applied to panoramas, and the resolution may be limited. On the other hand, animating perspective images individually can lead to inconsistencies between overlapping views.
@@ 4.3 Ablation Studies
We conduct ablation studies for both the animating and lifting phases of our methodology. In the animating phase, we highlight the importance of our spherical denoise strategy by replacing it with two basic animation techniques. In the lifting phase, we analyze the impact of excluding the Spatial-Temporal Geometry Alignment process and the temporal loss during the optimization of 4D representations.
@@
To animate the panorama into a panoramic video, a straightforward approach is to apply animators directly to the entire panorama. However, we observed that this strategy often results in minor motion, as shown in Fig. 5 (b) and Tab. 2 (Animate Pano.). This issue arises due to two main reasons: (1) animators are typically trained on perspective images with a narrow field of view (FoV), whereas panoramas have a 360 FoV with specific distortions under the equirectangular projection; (2) our panorama is high-resolution (4K), which exceeds the training distribution of most 2D animators and can easily cause out-of-memory issues, even with an 80GB VRAM graphics card. Thus the panoramas have to be down-sampled to a lower resolution (2K), causing a loss of details. Thus, we seek to animate on perspective views. Applying the animator to perspective views offers benefits such as reduced distortion and domain-appropriate input for the animator, allowing for smooth animation of high-resolution panoramas. However, animating perspective images separately can introduce inconsistencies between overlapping perspective views, as illustrated in Fig. 5 (c) and Tab. 2 (Animate Pers.). To resolve this challenge, we propose simultaneously denoising all perspective views and fusing them at each denoising step, in the spherical latent spaace, which capitalizes on the benefits of animating perspective views while ensuring cross-view consistency. The results are displayed in Fig. 5 (a) and Tab. 2 (Ours).
@@
Table 2: Different Animation Strategies in the Animating Phase. While the three strategies achieves similar visual quality (Q-Align), animating the entire panorama results in minor motion and reduced resolution (first row). Conversely, animating from perspective views leads to inconsistencies across different views (second row). This is supported by the “user agreement (UA)” study, where 70% of participants identified our method as view-consistent, in contrast to only 33% who consider animation from perspective views to be view-consistent.
@@
Animate Pano. $2048\times 1024$ 0.82 0.30 -
Animate Pers. $4096\times 2048$ 0.64 0.91 33%
Ours $4096\times 2048$ 0.85 1.27 70%
5 Conclusion
@@ Conclusion.
Appendix A Appendix
Due to space constraints in the main draft, we include supplementary details and experimental results in the appendix. Specifically, in Sec. B , we provide details about the acquisition process for the static panoramas used in our experiments. In Sec. C, we offer further explanation of the implementation for both the animation and lifting phases. Finally, in Sec. D, we describe the experimental setup and present additional results.
Appendix B Acquisition of Panoramas
The static panoramas used in the dataset of the main draft are generated by a text-to-panorama diffusion model, fine-tuned from stable diffusion  (Rombach et al., 2021) on SUN360. Similar to  (Feng et al., 2023), this model follows three steps: circular blending, super-resolution, and refinement. The panoramas are initially at a resolution of $6144\times 3072$ and then down-sampled to $4096\times 2048$ using the bi-linear interpolation.
Appendix C Implementation Details
In this section, we introduce the implementation details of the panoramic animator and the 4D lifting procedure.
Implementation of Spherical Representing
For the spherical representation, the continuous spherical mapping $\mathcal{S}_{I}:\mathbb{S}^{2}\rightarrow{}^{C}$ is instantiate as discrete point set $\mathcal{P}=\{p_{i}\}$, which is uniformly sampled from the sphere $\mathcal{S}_{I}$. We first initialize a icosahedron with 20 triangle faces $\{f_{i}|i=1,\cdots,20\}$ to approximate a real sphere $\mathbb{S}^{2}$. Then we uniformly sample a point set $P_{i}$ on each face $f_{i}$ and union all the point sets together as $\hat{\mathcal{P}}=\cup_{i=1}^{20}P_{i}$. We then obtain the discrete point set $\mathcal{P}$ by projecting $\hat{\mathcal{P}}$ onto the sphere $\mathbb{S}^{2}$ by $\mathcal{P}=\{{p_{i}}/{\|p_{i}\|}~{}|~{}p_{i}\in\hat{\mathcal{P}}\}$.
Figure 7: Visualizations: We provide more visual results. For each shown case we provide the input panorama, corresponding text prompt, and the rendering from two perspective views.
Panoramic Animation Phase
For the Panoramic Animator, we set the video length $L=14$, the channel number $c=9$, the latent code size $(h,w)=\frac{1}{8}(H,W)$, the perspective image size $p_{H}=p_{W}=\frac{1}{4}W$. The sphere is uniformly divided into 20 perspective views, each with 80 FOV. For the denoiser, the max denoising step is $25$. For the continuous optimization in Eq. 3, we use a close form, where each latent vector at each point on the sphere is the average of the latent vectors of the corresponding pixel on the perspective views that overlap it. The perspective denoiser is initiated as Animate-Anything Dai et al. (2023). The masks required by the denoiser are given by bounding boxes defined by user clicks.
Dynamic Panoramic Lifting Phase
In the lifting phase, similar to the animation phase, we choose the perspective view number $n=20$, each with 80 FOV. Each perspective view has a square shape, $P_{H}=P_{W}=\frac{1}{4}W$, where $W$ is the width of the original static panorama. In the Spatial-Temporal Geometric Alignment stage, the depth estimator $\Theta$ is implemented using MiDaS  (Ranftl et al., 2021; Birkl et al., 2023). The depth map from the perspective image is scaled according to the projection of the unit-length ray direction onto the camera orientation $\bm{d}$. Formally, if the pre-scaled depth is $d$ at point $p\in\hat{\mathcal{P}}$ introduced above, the scaled depth should be $d/\|p\|$. Additionally, for scenes without distinct boundaries, such as the sky, depth values for distant elements are assigned a finite value to support optimization.
Optimization Details
The hyper-parameters for optimization are set as follows: $\lambda_{\rm{depth}}=1,\lambda_{\rm{scale}}=0.1,\lambda_{\rm{shift}}=0.01$. We conduct Spatial-Temporal Geometry Alignment optimization over 3000 iterations, with $\lambda_{\rm{scale}}$ and $\lambda_{\rm{shift}}$ set to zero during the first 1500 iterations. For the 4D representation training stage, Gaussian parameters are optimized over 10000 iterations for each time stamp $t$. The hyper-parameters for this stage are defined as $\lambda_{\rm{rgb}}=1,\lambda_{\rm{temporal}}=\lambda_{\rm{sem}}=\lambda_{\rm{geo}}=0.05$, and the disturbance vector range $\alpha$ is varied at $0.05$, $0.1$, and $0.2$ during the $5400$, $6600$, and $9000$ iterations, respectively.
Appendix D Experimental Details
D.1 User Study Details
We conducted two user studies, gathering a total of 84 questionnaires from 42 users. For the “Quality (UC)”column in Tab. 1, we collected 42 questionnaires, each containing eight questions. Each question asked users to choose the bests video in term of visual quality from the perspective videos provided by different models. The user choice (UC) score of a method is the percentage of times the method’s video was selected as the best one, out of a total of 336 questions. Thus, the UC scores for all methods sum to 100%. For the “View-Consistency (UA)” column in Tab. 2, we collected another 42 questionnaires, with each questionnaire containing eight questions. Each question presented two videos from different views, both generated by the same method, and users were asked to determine whether the two videos were view-consistent. The user agreement (UA) score is the percentage of video pairs marked as view-consistent out of all the video pairs generated by the method. The UA scores do not necessarily sum to 100%. In the UC column of Tab. 1, among the 336 questions, users selected 4K4DGen 272 times, 3D-Cin. (circle) 40 times, and 3D-Cin. (zoomin) 24 times. In the UA column of Tab. 2, 118 out of 168 video pairs generated by “Our” were marked as consistent, while 56 out of 168 pairs from “Animate Pers” were considered consistent.
D.2 More Results
We provide additional qualitative results in Fig. 7. Furthermore, we highly recommend viewing the video renderings of 4K4DGen and comparisons to baseline models in the supplementary static HTML page for a more comprehensive and immersive experience.
Appendix E Ethics and Reproducibility Statement
Ethics Statement.
@@ Appendix E Ethics and Reproducibility Statement
Reproducibility.
We provide sufficient implementation details to reproduce our methodology in Sec. C, including the details of spherical denoiser, panoramic animator, dynamic panoramic lifting, etc. Furthermore, we will release the code in the future.